# Generative Recommender with End-to-End Learnable Item Tokenization

### v2

@@ Abstract.
Generative recommendation systems have gained increasing attention as an innovative approach that directly generates item identifiers for recommendation tasks. Despite their potential, a major challenge is the effective construction of item identifiers that align well with recommender systems. Current approaches often treat item tokenization and generative recommendation training as separate processes, which can lead to suboptimal performance. To overcome this issue, we introduce ETEGRec, a novel End-To-End Generative Recommender that unifies item tokenization and generative recommendation into a cohesive framework. Built on a dual encoder-decoder architecture, ETEGRec consists of an item tokenizer and a generative recommender. To enable synergistic interaction between these components, we propose a recommendation-oriented alignment strategy, which includes two key optimization objectives: sequence-item alignment and preference-semantic alignment. These objectives tightly couple the learning processes of the item tokenizer and the generative recommender, fostering mutual enhancement. Additionally, we develop an alternating optimization technique to ensure stable and efficient end-to-end training of the entire framework. Extensive experiments demonstrate the superior performance of our approach compared to traditional sequential recommendation models and existing generative recommendation baselines.
Generative Recommendation, Item Tokenization
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italy.booktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italyisbn: 979-8-4007-1592-1/25/07doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Recommendation systems
@@ 1. Introduction
Recently, drawing from the promising potential of generative language models (Sun et al., 2023; Yang et al., 2023), several studies have emerged to apply the generative paradigm in recommender systems (Rajput et al., 2023; Wang et al., 2024c; Tan et al., 2024; Wang et al., 2024a). Different from discriminative methods, the generative approach formulates the sequential recommendation task as a sequence-to-sequence problem and autoregressively generates the identifiers of target items. Specifically, it generally involves two main aspects, namely item tokenization and autoregressive generation, for developing the entire recommendation framework. For item tokenization, it basically refers to assigning a list of meaningful IDs for indexing or representing an item. Existing efforts include parameter-free methods based on the co-occurrence matrix (Petrov and Macdonald, 2023; Hua et al., 2023), text feature identifier (Li et al., 2023b; Di Palma, 2023; Tan et al., 2024), hierarchical clustering (Si et al., 2023; Wang et al., 2024c), and multi-level vector quantization (VQ) (Rajput et al., 2023; Qu et al., 2024; Wang et al., 2024a; Liu et al., 2024b). Furthermore, several recent studies attempt to improve the quality of item identifiers by introducing collaborative signals (Wang et al., 2024a), diversified regularization (Wang et al., 2024a), or multi-behavior information (Wang et al., 2024c; Liu et al., 2024a). For autoregressive generation, the encoder-decoder architecture (e.g., T5 (Raffel et al., 2020)) is the most widely used backbone due to its excellent capabilities in sequence modeling and generation. In addition, there are also some studies that aim to improve performance by adjusting the backbone architecture (Si et al., 2023; Wang et al., 2024c) or the learning objectives (Si et al., 2023).
@@ 2. Related Work
Sequential Recommendation. Sequential recommendation aims to predict the next item a user may interact with based on the user’s historical behavior sequences. Early studies (Rendle et al., 2010) primarily adhere to the Markov Chain assumption and focus on estimating the transition matrix. With the development of neural networks, various model architectures, such as Recurrent Neural Networks (RNN) (Hidasi et al., 2016; Tan et al., 2016), Convolutional Neural Networks (CNN) (Tang and Wang, 2018) and Graph Neural Networks (GNN) (Chang et al., 2021; Wu et al., 2019), are applied for sequential recommendation. Recently, Transformer (Vaswani et al., 2017)-based recommendation models (Kang and McAuley, 2018; Sun et al., 2019; Hao et al., 2023; Zhou et al., 2020) have achieved great success for effective sequential user modeling. SASRec (Kang and McAuley, 2018) utilizes Transformer decoder with unidirectional self-attention to capture user preference. BERT4Rec (Sun et al., 2019) proposes to encode the sequence by bidirectional attention and adopts the mask prediction task for training. S3Rec (Zhou et al., 2020) explores using the intrinsic data correlation as supervised signals to pre-train the sequential model for better user and item representations. Furthermore, several works exploit the abundant textual features of users and items to  (Zhang et al., 2019; Xie et al., 2022) enrich the user and item representations. In this work, we focus on exploring the generative paradigm for sequential recommendation.
@@
Generative Recommendation. Nowadays, generative recommendation has emerged as a next-generation paradigm for recommendation systems. In such a generative paradigm, the item sequence is tokenized into a token sequence and then fed into generative models to predict the tokens of the target item autoregressively. Generally, the generative paradigm can be considered as two main processes, i.e., item tokenization and generative recommendation. Existing approaches for item tokenization can be broadly categorized into parameter-free methods (Petrov and Macdonald, 2023; Hua et al., 2023; Yue et al., 2023; Tan et al., 2024; Si et al., 2023; Wang et al., 2024c), and deep learning methods based on multi-level vector quantization (VQ) (Rajput et al., 2023; Wang et al., 2024b; Qu et al., 2024; Wang et al., 2024a). For parameter-free methods, some studies, e.g., CID (Hua et al., 2023) and GPTRec (Petrov and Macdonald, 2023), apply matrix factorization to the co-occurrence matrix to derive item identifiers. Other works like SEATER (Si et al., 2023) and EAGER (Wang et al., 2024c) employ clustering of item embeddings to construct identifiers hierarchically. In addition, there are also some attempts (Li et al., 2023b; Di Palma, 2023; Harte et al., 2023; Yue et al., 2023; Tan et al., 2024) to use the textual metadata attached to items, e.g., titles and descriptions, as identifiers. While these non-parametric methods are highly efficient, they often suffer from length bias and fail to capture deeper collaborative relationships among items. Deep learning methods based on multi-level VQ instead develop more expressive item identifiers with equal length via the Deep Neural Networks (DNN). For instance, TIGER (Rajput et al., 2023) uses RQ-VAE to learn the codebooks. LETTER (Wang et al., 2024a) proposes to align quantized embeddings in RQ-VAE with collaborative embeddings to leverage both collaborative and semantic information.
@@ 3.1. Problem Formulation
As the task setting, following prior studies (Kang and McAuley, 2018; Sun et al., 2019), we consider a typical sequential recommendation scenario. Given the item set $\mathcal{I}$ and an interaction sequence $S=[i_{1},i_{2},\dots,i_{t}]$ from a target user $u$, sequential recommendation aims to predict the next item $i_{t+1}\in\mathcal{I}$ that user $u$ is likely to interact with. To approach this task, we take the generative paradigm by casting sequential recommendation as token sequence generation (Rajput et al., 2023): each item is indexed or represented by some ID identifier, and the task aims to generate the ID identifier of the future interacted item. As a straightforward approach, we can represent each item with its original item ID (e.g., a unique product number or a randomly assigned number). However, such an approach cannot effectively share similar semantics among different items, while often leading to a large item vocabulary.
In a more generic way, each item $i$ is represented by multiple tokens $[c_{1},\dots,c_{L}]$, where $L$ denotes the identifier length. In practice, $L$ can vary for different items, while we follow RQ-VAE (Zeghidour et al., 2022) to use the same identifier length for all the items, to reduce the potential length bias in item prediction (Wang et al., 2024a; Si et al., 2023). Following the convention in natural language processing (Raffel et al., 2020), we refer to the process of mapping an item into multiple tokens as item tokenization. In this representation scheme, the input interaction sequence $S$ can be first tokenized into the token sequence $X=[c^{1}_{1},c^{1}_{2},\dots,c^{t}_{L-1},c^{t}_{L}]$, where each item in $S$ is represented by its identifier (i.e., $L$ tokens) and $t$ is the original item index in the interaction sequence $S$. To obtain the item identifiers, previous studies either adopt the heuristic method (Li et al., 2023b; Geng et al., 2022; Hua et al., 2023) or employ a pre-learned tokenizer (Rajput et al., 2023; Wang et al., 2024a) for item tokenization. In this work, we consider devising an end-to-end approach for learning both the item tokenizer and the recommender backbone. The objective of generative recommendation is to first derive the token sequence $X$ and then generate the corresponding identifier of the target item $Y=[c^{t+1}_{1},\dots,c^{t+1}_{L}]$ at the $(t+1)$-th step. Formally, this task can be formulated into a typical sequence-to-sequence learning problem as follows:
(1) $P(Y|X)=\prod_{l=1}^{L}P(c^{t+1}_{l}|X,c^{t+1}_{1},..,c^{t+1}_{l-1}).$
@@ 3.2.1. Item Tokenizer
(10) $\mathcal{L}_{\operatorname{RQ}}$ $=\sum_{i=1}^{L}||\operatorname{sg}[\bm{v}_{l}]-\bm{e}^{l}_{c_{l}}||^{2}+\beta||\bm{v}_{l}-\operatorname{sg}[\bm{e}^{l}_{c_{l}}]||^{2},$
@@ 3.2.2. Generative Recommender
Token-level Seq2Seq Formulation. During training, the item-level user interaction sequence $S$ and the target item $i_{t+1}$ are first tokenized into the token-level sequence $X=[c^{1}_{1},c^{1}_{2},\dots,c^{t}_{1},c^{t}_{L}]$ and $Y=[c^{t+1}_{1},\dots,c^{t+1}_{L}]$ by $\mathcal{T}$. Then the corresponding token embeddings $\bm{E}^{X}\in\mathbb{R}^{|X|\times d_{h}}$ are fed into the generative recommender for user preference modeling, where $d_{h}$ is the hidden size of the recommender. Formally, we begin with token sequence encoding:
@@ 3.3. Recommendation-oriented Alignment
In previous work (Petrov and Macdonald, 2023; Si et al., 2023; Rajput et al., 2023; Wang et al., 2024a), the item tokenizer and the generative recommender are treated as two separate components: the tokenizer is often trained in the preprocessing stage to generate tokens for each item but is subsequently fixed during recommender training. Such an approach neglects the effect of item tokenization on the generative recommendation, which cannot adaptively learn more suitable tokenizers for the corresponding recommender. To address this limitation, an ideal approach is to jointly learn both the item tokenizer and the generative recommender for mutual enhancement between the two components. For this purpose, we devise two new training strategies for aligning the two components, namely sequence-item alignment and preference-semantic alignment.
@@ 3.3.2. Preference-Semantic Alignment
Through the above two alignment strategies, we can effectively enhance the associations between the two components during model optimization and thus can facilitate mutual enhancement by making necessary adaptations to each other.
@@ 3.4. Alternating Optimization
In general, we divide the training process into multiple cycles, each consisting of a fixed number of epochs. In the first epoch of each cycle, we optimize the item tokenizer based on Eq. (17) to improve the quality of item representations by the generative recommender. As for the rest epochs of each cycle, the item tokenizer is frozen and item tokens remain fixed during the generative recommender training process. This approach ensures stable optimization when conducting the recommendation-oriented alignment.
@@ 4. Experiments
4.1. Experiment Setup
4.1.1. Dataset
We conduct experiments on three subsets of the most recent Amazon 2023 review data (Hou et al., 2024) to evaluate our approach, including “Musical Instruments”, “Video Games”, and “Industrial Scientific”. All these datasets comprise user review data from May 1996 to September 2023. Following previous works (Zhou et al., 2020; Zheng et al., 2023), we apply the 5-core filter to exclude unpopular users and items with less than five interaction records. Then, we construct user behavior sequences according to the chronological order and uniformly set the maximum item sequence length to 50. The statistics of preprocessed datasets are shown in Table 2.
@@ 4.1.2. Baseline Models
The baseline models we adopt for comparison include the following two categories:
(1) Traditional sequential recommendation models:
  • Caser (Tang and Wang, 2018) utilizes horizontal and vertical convolutional filters to model user behavior sequences.
  • HGN (Ma et al., 2019) employs hierarchical gating networks to capture both long-term and short-term user interests from item sequences.
  • GRU4Rec (Hidasi et al., 2016) is an RNN-based sequential recommender that uses GRU for user behavior modeling.
  • BERT4Rec (Sun et al., 2019) introduces bidirectional Transformer and mask prediction tasks into sequential recommendation for user preference modeling.
  • SASRec (Kang and McAuley, 2018) adopts the unidirectional Transformer to model user behaviors and predict the next item.
  • FMLP-Rec (Zhou et al., 2022) proposes an all-MLP sequential recommender with learnable filters, which can effectively reduce user behavior noise.
  • FDSA (Zhang et al., 2019) emphasizes the transformation patterns between item features by separately modeling both item-level and feature-level sequences using self-attention networks.
  • S3-Rec (Zhou et al., 2020) incorporates mutual information maximization into sequential recommendation for model pre-training, learning the correlation between items and attributes to improve recommendation performance.
(2) Generative recommendation models:
  • SID (Hua et al., 2023) sequentially encodes item IDs according to user interaction order and utilizes them as item identifiers for generative recommendation.
  • CID (Hua et al., 2023) integrates collaborative knowledge into LLM-based generative recommender by generating item identifiers through spectral clustering on item co-occurrence graphs.
  • TIGER (Rajput et al., 2023) leverages text embedding to construct semantic IDs for items and adopts the generative retrieval paradigm for sequential recommendation.
  • TIGER-SAS (Rajput et al., 2023) uses the item embeddings from trained SASRec instead of text embeddings to construct semantic IDs, which enables item identifiers to imply collaborative prior knowledge.
  • LETTER (Wang et al., 2024a) designs a learnable tokenizer by integrating hierarchical semantics, collaborative signals, and code assignment diversity.
4.1.3. Evaluation Settings
4.1.4. Semantic ID Generation
4.1.5. Implementation Details
Traditional recommendation baselines are implemented by an open-source recommendation framework RecBole (Zhao et al., 2022). For CID, SID [^2] https://github.com/Wenyueh/LLM-RecSys-ID and LETTER [^3] https://github.com/HonghuiBao2000/LETTER, we utilize their official implementations. For TIGER and TIGER-SAS, we follow the implementation details provided in the original paper (Rajput et al., 2023). To ensure fair comparisons, the item embedding dimension is set to 128 for all models, except for SID and CID, which retain their default dimension of 768.

### v3

@@ Abstract.
Generative recommender systems have gained increasing attention as an innovative approach that directly generates item identifiers for recommendation tasks. Despite their potential, a major challenge is the effective construction of item identifiers that align well with recommender systems. Current approaches often treat item tokenization and generative recommendation training as separate processes, which can lead to suboptimal performance. To overcome this issue, we introduce ETEGRec, a novel End-To-End Generative Recommender that unifies item tokenization and generative recommendation into a cohesive framework. Built on a dual encoder-decoder architecture, ETEGRec consists of an item tokenizer and a generative recommender. To enable synergistic interaction between these components, we propose a recommendation-oriented alignment strategy, which includes two key optimization objectives: sequence-item alignment and preference-semantic alignment. These objectives tightly couple the learning processes of the item tokenizer and the generative recommender, fostering mutual enhancement. Additionally, we develop an alternating optimization technique to ensure stable and efficient end-to-end training of the entire framework. Extensive experiments demonstrate the superior performance of our approach compared to traditional sequential recommendation models and existing generative recommendation baselines. Our code is available at https://github.com/RUCAIBox/ETEGRec.
Generative Recommendation; Item Tokenization
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729989isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
@@ 1. Introduction
Recently, drawing from the promising potential of generative language models (Sun et al., 2023; Yang et al., 2023), several studies have emerged to apply the generative paradigm in recommender systems (Rajput et al., 2023; Wang et al., 2024c; Tan et al., 2024; Wang et al., 2024a). Different from discriminative methods, the generative approach formulates the sequential recommendation task as a sequence-to-sequence problem and autoregressively generates the identifiers of target items. Specifically, it generally involves two main aspects, namely item tokenization and autoregressive generation, for developing the entire recommendation framework. For item tokenization, it basically refers to assigning a list of meaningful IDs for indexing or representing an item. Existing efforts include parameter-free methods based on the co-occurrence matrix (Petrov and Macdonald, 2023; Hua et al., 2023), text feature identifier (Li et al., 2023b; Di Palma, 2023; Tan et al., 2024), hierarchical clustering (Si et al., 2024; Wang et al., 2024c), and multi-level vector quantization (VQ) (Rajput et al., 2023; Qu et al., 2024; Wang et al., 2024a; Liu et al., 2024b). Furthermore, several recent studies attempt to improve the quality of item identifiers by introducing collaborative signals (Wang et al., 2024a), diversified regularization (Wang et al., 2024a), or multi-behavior information (Wang et al., 2024c; Liu et al., 2024a). For autoregressive generation, the encoder-decoder architecture (e.g., T5 (Raffel et al., 2020)) is the most widely used backbone due to its excellent capabilities in sequence modeling and generation. In addition, there are also some studies that aim to improve performance by adjusting the backbone architecture (Si et al., 2024; Wang et al., 2024c) or the learning objectives (Si et al., 2024).
@@ 2. Related Work
Sequential Recommendation. Sequential recommendation aims to predict the next item a user may interact with based on the user’s historical behavior sequences. Early studies (Rendle et al., 2010) primarily adhere to the Markov Chain assumption and focus on estimating the transition matrix. With the development of neural networks, various model architectures, such as Recurrent Neural Networks (RNN) (Hidasi et al., 2016; Tan et al., 2016), Convolutional Neural Networks (CNN) (Tang and Wang, 2018) and Graph Neural Networks (GNN) (Chang et al., 2021; Wu et al., 2019), are applied for sequential recommendation. Recently, Transformer (Vaswani et al., 2017)-based recommendation models (Kang and McAuley, 2018; Sun et al., 2019; Hao et al., 2023; Zhou et al., 2020) have achieved great success for effective sequential user modeling. SASRec (Kang and McAuley, 2018) utilizes Transformer decoder with unidirectional self-attention to capture user preference. BERT4Rec (Sun et al., 2019) proposes to encode the sequence by bidirectional attention and adopts the mask prediction task for training. S3Rec (Zhou et al., 2020) explores using the intrinsic data correlation as supervised signals to pre-train the sequential model for better user and item representations. Furthermore, several works (Zhang et al., 2019; Xie et al., 2022) exploit the abundant textual features of users and items to enrich the user and item representations. In this work, we focus on exploring the generative paradigm for sequential recommendation.
@@
Generative Recommendation. Nowadays, generative recommendation has emerged as a next-generation paradigm for recommendation systems. In such a generative paradigm, the item sequence is tokenized into a token sequence and then fed into generative models to predict the tokens of the target item autoregressively. Generally, the generative paradigm can be considered as two main processes, i.e., item tokenization and generative recommendation. Existing approaches for item tokenization can be broadly categorized into parameter-free methods (Petrov and Macdonald, 2023; Hua et al., 2023; Yue et al., 2023; Tan et al., 2024; Si et al., 2024; Wang et al., 2024c), and deep learning methods based on multi-level vector quantization (VQ) (Rajput et al., 2023; Wang et al., 2024b; Qu et al., 2024; Wang et al., 2024a). For parameter-free methods, some studies, e.g., CID (Hua et al., 2023) and GPTRec (Petrov and Macdonald, 2023), apply matrix factorization to the co-occurrence matrix to derive item identifiers. Other works like SEATER (Si et al., 2024) and EAGER (Wang et al., 2024c) employ clustering of item embeddings to construct identifiers hierarchically. In addition, there are also some attempts (Li et al., 2023b; Di Palma, 2023; Harte et al., 2023; Yue et al., 2023; Tan et al., 2024) to use the textual metadata attached to items, e.g., titles and descriptions, as identifiers. While these non-parametric methods are highly efficient, they often suffer from length bias and fail to capture deeper collaborative relationships among items. Deep learning methods based on multi-level VQ instead develop more expressive item identifiers with equal length via the Deep Neural Networks (DNN). For instance, TIGER (Rajput et al., 2023) uses RQ-VAE to learn the codebooks. LETTER (Wang et al., 2024a) proposes to align quantized embeddings in RQ-VAE with collaborative embeddings to leverage both collaborative and semantic information.
@@ 3.1. Problem Formulation
As the task setting, following prior studies (Kang and McAuley, 2018; Sun et al., 2019), we consider a typical sequential recommendation scenario. Given the item set $\mathcal{I}$ and an interaction sequence $S=[i_{1},i_{2},\dots,i_{t}]$ from a target user $u$, sequential recommendation aims to predict the next item $i_{t+1}\in\mathcal{I}$ that user $u$ is likely to interact with. To approach this task, we take the generative paradigm by casting sequential recommendation as token sequence generation (Rajput et al., 2023): each item is indexed or represented by a specific ID identifier, and the task aims to generate the ID identifier of the future interacted item. As a straightforward approach, we can represent each item with its original item ID (e.g., a unique product number or a randomly assigned number). However, such an approach cannot effectively share similar semantics among different items, while often leads to a large item vocabulary.
In a more generic way, each item $i$ is represented by multiple tokens $[c_{1},\dots,c_{L}]$, where $L$ denotes the identifier length. In practice, $L$ can vary for different items, while we follow RQ-VAE (Zeghidour et al., 2022) to use the same identifier length for all the items, to reduce the potential length bias in item prediction (Wang et al., 2024a; Si et al., 2024). Following the convention in natural language processing (Raffel et al., 2020), we refer to the process of mapping an item into multiple tokens as item tokenization. In this representation scheme, the input interaction sequence $S$ can be first tokenized into the token sequence $X=[c^{1}_{1},c^{1}_{2},\dots,c^{t}_{L-1},c^{t}_{L}]$, where each item in $S$ is represented by its identifier (i.e., $L$ tokens) and $t$ is the original item index in the interaction sequence $S$. To obtain the item identifiers, previous studies either adopt the heuristic method (Li et al., 2023b; Geng et al., 2022; Hua et al., 2023) or employ a pre-learned tokenizer (Rajput et al., 2023; Wang et al., 2024a) for item tokenization. In this work, we consider devising an end-to-end approach for learning both the item tokenizer and the recommender backbone. The objective of generative recommendation is to first derive the token sequence $X$ and then generate the corresponding identifier of the target item $Y=[c^{t+1}_{1},\dots,c^{t+1}_{L}]$ at the $(t+1)$-th step. Formally, this task can be formulated into a typical sequence-to-sequence learning problem as follows:
(1) $P(Y|X)=\prod_{l=1}^{L}P(c^{t+1}_{l}|X,c^{t+1}_{1},..,c^{t+1}_{l}).$
@@ 3.2.1. Item Tokenizer
(10) $\mathcal{L}_{\operatorname{RQ}}$ $=\sum_{l=1}^{L}||\operatorname{sg}[\bm{v}_{l}]-\bm{e}^{l}_{c_{l}}||^{2}+\beta||\bm{v}_{l}-\operatorname{sg}[\bm{e}^{l}_{c_{l}}]||^{2},$
@@ 3.2.2. Generative Recommender
Token-level Seq2Seq Formulation. During training, the item-level user interaction sequence $S$ and the target item $i_{t+1}$ are first tokenized into the token-level sequence $X=[c^{1}_{1},c^{1}_{2},\dots,c^{t}_{L-1},c^{t}_{L}]$ and $Y=[c^{t+1}_{1},\dots,c^{t+1}_{L}]$ by $\mathcal{T}$. Then the corresponding token embeddings $\bm{E}^{X}\in\mathbb{R}^{|X|\times d_{h}}$ are fed into the generative recommender for user preference modeling, where $d_{h}$ is the hidden size of the recommender. Formally, we begin with token sequence encoding:
@@ 3.3. Recommendation-oriented Alignment
In previous work (Petrov and Macdonald, 2023; Si et al., 2024; Rajput et al., 2023; Wang et al., 2024a), the item tokenizer and the generative recommender are treated as two separate components: the tokenizer is often trained in the preprocessing stage to generate tokens for each item but is subsequently fixed during recommender training. Such an approach neglects the effect of item tokenization on the generative recommendation, which cannot adaptively learn more suitable tokenizers for the corresponding recommender. To address this limitation, an ideal approach is to jointly learn both the item tokenizer and the generative recommender for mutual enhancement between the two components. For this purpose, we devise two new training strategies for aligning the two components, namely sequence-item alignment and preference-semantic alignment.
@@ 3.3.2. Preference-Semantic Alignment
Through the above two alignment strategies, we can effectively enhance the association between the two components during model optimization and thus can facilitate mutual enhancement by making necessary adaptations to each other.
@@ 3.4. Alternating Optimization
In general, we divide the training process into multiple cycles, each consisting of a fixed number of epochs. In the first epoch of each cycle, we optimize the item tokenizer based on Eq. (17) to improve the quality of item representations by the generative recommender. As for the rest epochs of each cycle, the item tokenizer is frozen and item tokens remain fixed during the generative recommender training process. This alternation continues until the item tokenizer converges, after which we permanently freeze it and fully train the generative recommender to convergence This approach ensures stable optimization when conducting the recommendation-oriented alignment.
@@ 4. Experiments
4.1. Experiment Setup
4.1.1. Dataset
We conduct experiments on three subsets of the most recent Amazon 2023 review data (Hou et al., 2024) to evaluate our approach, including “Musical Instruments”, “Video Games”, and “Industrial Scientific”. All these datasets comprise user review data from May 1996 to September 2023. Following previous works (Zhou et al., 2020; Zheng et al., 2024), we apply the 5-core filter to exclude unpopular users and items with less than five interaction records. Then, we construct user behavior sequences according to the chronological order and uniformly set the maximum item sequence length to 50. The statistics of preprocessed datasets are shown in Table 2.
@@ 4.1.2. Baseline Models
The baseline models we adopt for comparison include the following two categories: (1) Traditional sequential recommendation models: Caser (Tang and Wang, 2018) utilizes horizontal and vertical convolutional filters to model user behavior sequences. HGN (Ma et al., 2019) employs hierarchical gating networks to capture both long-term and short-term user interests from item sequences. GRU4Rec (Hidasi et al., 2016) is an RNN-based sequential recommender that uses GRU for user behavior modeling. BERT4Rec (Sun et al., 2019) introduces bidirectional Transformer and mask prediction tasks into sequential recommendation for user preference modeling. SASRec (Kang and McAuley, 2018) adopts the unidirectional Transformer to model user behaviors and predict the next item. FMLP-Rec (Zhou et al., 2022) proposes an all-MLP sequential recommender with learnable filters, which can effectively reduce user behavior noise. FDSA (Zhang et al., 2019) emphasizes the transformation patterns between item features by separately modeling both item-level and feature-level sequences using self-attention networks. S3-Rec (Zhou et al., 2020) incorporates mutual information maximization into sequential recommendation for model pre-training, learning the correlation between items and attributes to improve recommendation performance. (2) Generative recommendation models: SID (Hua et al., 2023) sequentially encodes item IDs according to user interaction order and utilizes them as item identifiers for generative recommendation. CID (Hua et al., 2023) integrates collaborative knowledge into LLM-based generative recommender by generating item identifiers through spectral clustering on item co-occurrence graphs. TIGER (Rajput et al., 2023) leverages text embedding to construct semantic IDs for items and adopts the generative retrieval paradigm for sequential recommendation. TIGER-SAS (Rajput et al., 2023) uses the item embeddings from trained SASRec instead of text embeddings to construct semantic IDs, which enables item identifiers to imply collaborative prior knowledge. LETTER (Wang et al., 2024a) designs a learnable tokenizer by integrating hierarchical semantics, collaborative signals, and code assignment diversity.
@@
4.1.3. Evaluation Settings
4.1.4. Semantic ID Generation
4.1.5. Implementation Details
Traditional recommendation baselines are implemented by an open-source recommendation framework RecBole (Zhao et al., 2021, 2022). For CID, SID (Hua et al., 2023) and LETTER (Wang et al., 2024a), we utilize their official implementations. For TIGER and TIGER-SAS, we follow the implementation details provided in the original paper (Rajput et al., 2023). To ensure fair comparisons, the item embedding dimension is set to 128 for all models, except for SID and CID, which retain their default dimension of 768.